Fermi National Laboratory
 

Operation Summary for November 9, 2009 - November 11, 2009

 

Monday Nov. 9

The day shift began with Operations monitoring the Tevatron (TeV) store 7333 (169.96E30) and the Recycler stash (109.8E10), with the Antiproton Source stacking (11.3E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

During the day shift, RF personnel replaced a MI RF station (MIRF6) modulator.

During the evening shift, Operations terminated store 7333 at 6:33 PM and TeV studies followed. Operations put the TeV into HEP shot setup at 10 PM, but then the H- Source ground vacuum pump failed. The shot setup was abandoned and Operations called in H- Source personnel. They repaired the vacuum system and began recovering the source.

Tuesday Nov. 10

The midnight shift began with Operations monitoring the stash (371.0E10) and the stack (37.4E10), and with NuMI and MiniBooNE off due to H- Source repairs.

During the midnight shift, Operations reported at 1:35 AM that the H- Source was brought back online. Operations resumed sending beam to Pbar, NuMI, and MiniBooNE. The MI low-level RF (LLRF) system and dampers required a reset, which inhibited stacking and beam to NuMI for 45 minutes. Operations put the TeV into shot setup at 4 AM and established store 7340 at 5:35 AM with an average initial luminosity (AIL) of 278.0E30. A controls failure in Pbar occurred at 5:50 AM that inhibited stacking. Controls personnel came in and investigated a faulty broadcast module for the global clock system. Repairs lasted the rest of the midnight shift.

During the day shift, controls personnel reported finding a bad input for a repeater at AP0. They replaced the repeater, which resolved the Pbar clock problems. MIRF13 tripped off and caused the NuMI beam permit to trip off. NuMI reset. An RF technician investigated MIRF13. The Neutron Therapy Facility completed treatment on their last patient for the day at 12:40 PM.

During the evening shift, short downtimes were caused by BRF15 tripping off twice during the shift and MIRF2 once; operators bypassed MIRF2. Booster personnel tuned their beam for most of the shift; NuMI was set to 10 turns, which was a slight drop in their intensity. Operations terminated store 7340 at 11:26 PM and put the TeV into shot setup.

Wednesday Nov. 11

The midnight shift began with the TeV in shot setup, with Pbar stacking (41.9E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations established store 7341 at 1:07 AM with an AIL of 255.7E30. BRF19 tripped off, it reset. A MI Lambertson (LAM42) tripped off at 3:45 AM; it reset, but would not turn on. Operators called in EE Support personnel and loaded an energy conservation timeline into the Time Line Generator. EE Support personnel reported finding two failed SCRs and an LCW leak in the Lambertson power supply.

The Plans for Wednesday

The plans for today are to complete repairs on LAM42 and then resume stacking and stashing. The Run Coordinator will allow a TeV cryo system front-end reboot. He will also allow a MI-8 transfer line study and a Booster study.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/11/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory